2)     Izabela, a 10 year girl, asks for a red balloon for Christmas. Her father purchases a red balloon with ‘Merry Christmas Izzy’ written on it. Izabela blows the balloon into a 30.0 cm diameter balloon using helium gas. Assume the balloon is a sphere.

a.      Determine the buoyant force on the balloon.

b.      If the mass of the rubber balloon is 7.0 g, determine the net upward force on the balloon. Will the balloon rise?

 

You may assume that the density of helium is 1.28 kg/m3 and that .


1
Expert's answer
2021-11-30T09:50:00-0500

a. "F_B=\rho_a g V=\rho_a g (4/3)\pi r^3="


"=1.29\cdot9.8\cdot 4\cdot3.14\cdot0.15^3/3=0.18 \ (N)"


b. "F=F_B-m_gg-m_rg="


"=0.18-0.178\cdot 4\cdot 3.14\cdot0.15^3\cdot9.8/3-0.007\cdot9.8="


"\approx0.09\ (N)" . The balloon will rise.